# What is glacier sliding?

### v1

@@ 2 Inner-outer flows
The height at which inner-flow processes become negligible will also vary as a function of the spatial scale under consideration, heterogeneity in bed properties, and the roughness characteristics of the bed topography. For example, inner-flow processes will be important at a greater height above the bed if considering an area of rough topography around a planar soft-bed region, than if considering the planar soft-bed region in isolation. In most situations however, a reasonable norm is already defined for setting the position of the model bed surface. BedMachine (Morlighem and others,, 2017) and BedMap (Fretwell and others,, 2013) products are used for the Greenland and Antarctic ice sheets respectively while a more fragmented approach is taken for mountain glaciers (e.g. Welty and others,, 2020; Grab and others,, 2021). These bed topography models are significantly smoother than the roughness suggested by geostatistical analysis of radar flight lines (MacKie and others,, 2021; Law and others,, 2023) and deglaciated forelands immediately adjacent to the ice sheets, but function as convenient thresholds for separating inner and outer flows. Crucially, we note that the imperfect representation inherent to subglacial bed topography products means a degree of ice deformation will unavoidably already be incorporated within existing sliding parameterisations even if its inclusion has not been previously formalised. The error associated with a collapsed inner flow is therefore not introducing new errors, but rather providing a way to quantify and address existing ones.
Using this description of the inner flow, the net basal velocity at the model bed surface (m a-1) can then be treated as the sum of individual sliding processes up to $\mathcal{P}_{n}$ with
$u_{b}=\mathcal{P}_{1}+\mathcal{P}_{2}+...+\mathcal{P}_{n}.$ (2)
and the net basal traction (Pa) as
$\tau_{b}=\hat{\mathcal{P}_{1}}+\hat{\mathcal{P}_{2}}+...+\hat{\mathcal{P}_{n}}$ (3)
where hats ( $\hat{}$ ) refer to the resistance opposing each sliding process over the area considered. This is displayed schematically in Fig 6. Expanding Eqs. 2 and 3 to cover an area $R$, which may represent a model cell (e.g. Fig. 3), gives
$u_{b}\rvert_{R}=\frac{1}{R}\iint_{R}\mathcal{P}_{1}+\mathcal{P}_{2}+...+\mathcal{P}_{n}\,\mathrm{d}R$ (4)
and
$\tau_{b}\rvert_{R}=\frac{1}{R}\iint_{R}\hat{\mathcal{P}_{1}}+\hat{\mathcal{P}_{2}}+...+\hat{\mathcal{P}_{n}}\,\mathrm{d}R\,.$ (5)
This framework facilitates quantification of the role of each sub-process in the inner flow, though a full representation will be further complicated if sliding sub-processes are in fact represented as functions of one another (e.g. $\hat{\mathcal{P}_{1}}=f(\hat{\mathcal{P}_{2}},\hat{\mathcal{P}_{3}})$).
Figure 6: Schematics of inner flow velocity and traction components for planar soft-bed sliding. Sub-processes are sediment deformation ($\mathcal{P}_{1}$), slip at the ice-till interface ($\mathcal{P}_{2}$), and viscous deformation and regelation around clasts that are substantially larger than the surrounding till matrix ($\mathcal{P}_{3}$). This reflects the transition zone between viscous deformation and sediment shearing modes detailed in Zoet and Iverson, (2020). An example featuring more processes is Appendix C.
@@ 3.1 Sliding over ‘soft’ (sediment) beds
More recently, Zoet and Iverson, (2020) used a ring of temperate ice $\sim$15 cm thick between the rigid shear ring and underlying sediment. In this setting, rate-strengthening behaviour occurs at lower velocities (below around 50 m a-1 in sediment with clasts, Fig. 4a) as ice viscously deforms and regelates around static clasts before the clasts begin to plough through the finer-grained matrix at a stress limited by the till’s Coulomb strength (Fig. 4b). The net result of the Zoet and Iverson, (2020) experiment is a regularised-Coulomb relationship (Fig. 1, Table 1) where the sliding interface is defined up 15 cm above the ice-till interface. The relationship of Zoet and Iverson, (2020) to form drag as proposed by Minchew and Joughin, (2020) is covered in Section 3.3.
@@ 3.2 Sliding over ‘hard’ (bedrock) beds
The slip of ice over a planar surface (Fig. 4c) is closer to the assumptions of standard Coulomb behaviour (Coulomb,, 1785; Desplanques,, 2014), though plane-normal velocity components, englacial clasts, and possible clast-scale cavitation present complications (Hoffman and others,, 2022). Rather than a sliding relationship of the form of Eq. 1, basal traction in these studies is often reported as a static $\mu_{s}$ or transient $\mu_{t}$ coefficient of friction, $\mu_{s}\>\textup{or}\>\mu_{t}=\frac{\tau_{b}}{N}$ where $\mu_{s}$ describes the force required to move the plane from stationary and $\mu_{t}$ describes the force required to keep the plane in motion. Shear-ring studies using temperate ice and englacial sediment content of 0-20% over macroscopically flat granite find a $\mu_{s}$ of 0.02-0.05 (Barnes and others,, 1971; Zoet and others,, 2013; McCarthy and others,, 2017; Thompson,, 2020), lower than teflon-on-teflon with no lubrication (Fetfatsidis and others,, 2013). In the above studies, sliding may be stick-slip or steady-state (outlined further in Section 3.4) depending on water pressure and particular material properties. Transient coefficients of friction between 0.05-0.08 were obtained from subglacial access to the ice-rock interface of the Engabreen outlet glacier in Norway (Iverson and others,, 2003; Cohen and others,, 2005). Unexpectedly high shear traction values of up to 0.5 MPa accompany the low $\mu_{t}$ values at Engabreen and these are still not fully explained, suggesting a possible gap in understanding – see Thompson and others, (2020) for a more in depth discussion. The transient coefficient of friction is much higher for cold ice (0.5 at -20oC) making sub-temperate slip very low, but potentially non-negligible (Barnes and others,, 1971; McCarthy and others,, 2017; Atkins,, 2013; Mantelli and others,, 2019; Section 3.7).
@@
Sliding over rough hard beds has also received much attention. This began with the classical work of Weertman, (1957), who considered ice motion over isolated cuboids (akin to the geometry in Fig. 4e) with no ice-bed separation through two processes: regelation and enhanced creep (Fig. 4e). In regelation, ice moves without deformation. Temperate ice approaching an obstacle is subject to greater pressure, lowering its melting point. Consequently, some of the ice melts at the upstream face, and the meltwater is driven to the downstream face by the pressure gradient. Here, where the pressure is lower and the melting point is higher, the meltwater refreezes. The higher temperature at the downflow side of the obstacle results in heat flow back through the obstacle, completing the cycle. Regelation is an ongoing process in many subglacial settings (Kamb and Lachapelle,, 1964; Kamb,, 1970; Hallet and others,, 1978; Hubbard and Sharp,, 1993; Iverson,, 2000; Cook and others,, 2011; Rempel and Meyer,, 2019), including in soft bed settings outside the scope of the original Weertman, (1957) paper. However, the requirement for return heat flow causes regelation rate to vary inversely with obstacle wavelength, limiting the importance of regelation as a sliding process to obstacles smaller than a few cm Weertman, (1957). Regelation is therefore generally not viewed as a significant contributor to overall sliding rates (MacAyeal,, 2019) and has received comparatively little attention since the mid 1990s, though further contemporary study would not be unwelcome.
@@
Enhanced creep is prompted by stress concentrations at bedrock obstacles, leading to locally increased strain rates. Functionally, this means enhanced creep can be considered as a synonym for form drag (Section 3.3). Conversely to regelation, the contribution of enhanced creep to glacier sliding in Weertman’s 1957 study increases with obstacle size. Focusing on the obstacle sizes that produce the greatest rate of sliding when regelation and enhanced creep act in unison (millimetre- to metre-scale) Weertman, (1957) reaches a value of $m$ in a power law of $(n+1)/2=2$ with $C$ representing a combination of bed-geometry characteristics. Technically, values of $m$ deviating from 2 when $n=3$ are therefore not physically based on the original Weertman theory making ‘power-law’ or ‘Weertman-type’ sliding a more suitable descriptor. Sliding over a sinusoidal two-dimensional bed with no ice-bed separation and a linear ice rheology produces a linear sliding relationship ($m=1$ in Weertman-type sliding) (Nye,, 1969), as does Weertman sliding with $n=1$.
@@
Weertman-type sliding, usually with $m=3$, is the most commonly used sliding relationship in the latest round of ISMIP experiments (with linear sliding also frequently used) (Goelzer and others,, 2020; Seroussi and others,, 2020), yet the theoretical underpinnings behind these relationships do not hold up to scrutiny – including by Weertman’s own reappraisal (Weertman,, 1979). Weertman’s original theory features a non-linear rheology following Glen, (1955), but it is apparent that no glacier beds actually have the morphology used by Weertman, (1957), and even changes to a different unrealistic bed geometry produces a materially different relationship (Lliboutry,, 1979). Weertman sliding is further challenged by ubiquitous evidence and theoretical support for subglacial cavitation, or regions of bounded ponded water at high pressure between the glacier sole and underlying hard or soft bed, which form when the pressure on the downstream face of an obstacle falls below a critical level (Fig. 4g and e.g. Walder and Hallet,, 1979; Kamb,, 1987; Hooke,, 1989; Helanow and others,, 2021). Weertman sliding also does not provide an explanation for increased sliding rates when subglacial water pressure increases. The persistent utility of Weertman-type sliding in heuristic studies, though generally with $m\geq 3$, despite these major problems is covered in Sections 5 and 6.4.
@@
Subglacial cavities (Fig. 4g) act to reduce contact between the glacier sole and the ice bed. In the two-dimensional formulations of Lliboutry, (1968), Iken, (1981), Schoof, (2005), and Gagliardini and others, (2007), amongst others, where the bed is comprised of frictionless sinusoids, this produces an upper bound for basal traction, followed by rate-weakening behaviour (Fig. 1). Put mathematically, ’Iken’s bound’ as coined by Schoof, (2005) is given as
@@
where $\theta$ is the maximum up-slope angle between the bed and the mean flow direction of the ice (Iken,, 1981; Schoof,, 2005). Roldan-Blasco and others, (2022) suggest that in the simplified two-dimensional framework of compound sinusoids non-negligible friction at the ice-bed interface will not significantly alter the form of the sliding relationship or the existence of Iken’s bound. However, the geometry used in these studies can also be considered unrealistic, with the absence of a third dimension preventing lateral escape of subglacial water that would act to reduce the cavity size (Gimbert and others,, 2021). Helanow and others, (2019) and Helanow and others, (2021) address this and show that realistic three-dimensional bed topographies up to 25 m still result in bounded basal traction, but that significant rate-weakening behaviour is not expected (the regularised-Coulomb relationship in Fig. 1). The validity of the assumptions underlying Eq. 6 when considering inner flows over larger areas and rough topography is covered in Section 6.3.

### v3

@@ 2 Inner-outer flows
The height at which inner-flow processes become negligible will also vary as a function of the spatial scale under consideration, heterogeneity in bed properties, and the roughness characteristics of the bed topography. For example, inner-flow processes will be important at a greater height above the bed if considering an area of rough topography around a planar soft-bed region, than if considering the planar soft-bed region in isolation. In most situations however, a reasonable norm is already defined for setting the position of the model bed surface. BedMachine (Morlighem and others,, 2017) and BedMap (Fretwell and others,, 2013) products are used for the Greenland and Antarctic ice sheets respectively while a more fragmented approach is taken for mountain glaciers (e.g. Welty and others,, 2020; Grab and others,, 2021). These bed topography models are significantly smoother than the roughness suggested by geostatistical analysis of radar flight lines (MacKie and others,, 2021; Law and others,, 2023) and deglaciated forelands immediately adjacent to the ice sheets, but function as convenient thresholds for separating inner and outer flows. And, even if there is full knowledge of the bed topography (for example modelling the palaeo-Scandinavian Ice Sheet), model resolution dictates some inherent reduction in fidelity of bed topography. Crucially, we note that the imperfect representation inherent to subglacial bed topography products means a degree of ice deformation will unavoidably already be incorporated within existing sliding parameterisations even if its inclusion has not been previously formalised. The error associated with a collapsed inner flow is therefore not introducing new errors, but rather providing a way to quantify and address existing ones.
Using this description of the inner flow, the net basal traction, $\tau_{b}$ or $\boldsymbol{\tau}_{b}=(\tau_{b,x},\tau_{b,y})$ for higher-dimensional settings, at the model bed surface can then be treated as the sum of multiple individual sliding processes and upslope bed resistance. Here, we separate these inner-flow processes into traction components for the following categories: resistance to discrete slip at the ice-bed interface represented as $\mathcal{S}$, resistance to distributed viscous deformation within the ice as as $\mathcal{D}$, and geometric resistance to flow from up-slope obstacles as $\mathcal{G}$, with each printed in bold font to represent vector components. In a one dimensional case we can relate $\mathcal{S}$ to the slip velocity at the ice-bed interface, $u_{s}$, $\mathcal{D}$ to the velocity gradient between the ice-bed interface and the inner-outer boundary, $u$, and $\mathcal{G}$ to the slope orientation, $\theta$, though we remain agonistic about the exact functions tying these terms together for now. Summing these components for a simple one-dimensional case gives
$\tau_{b}=\sum_{i}\mathcal{S}_{i}(u_{s})+\sum_{i}\mathcal{D}_{i}(u)+\mathcal{G}_{i}(\theta)$ (2)
for each sub-process $i$ within each category.
Considering next a region, $R$, that could be taken as a model cell (Fig. 6), and which may contain sliding processes that vary through space, where the plane defined by $x$ and $y$ follows the average of the plane defining the inner-outer boundary over the region which is also parallel to $\boldsymbol{\tau}_{b}$ and $\boldsymbol{u}_{b}$, the vector defining velocity at the inner-outer boundary. We can then define each sub-process as additionally being a function of space using an indicator function $I_{R_{i}}(x,y)$ dependent on a sub-region $R_{i}$ where
$I_{R_{i}}(x,y)=\begin{cases}1&\text{if }(x,y)\in R_{i}\,,\\ 0&\text{otherwise}\,.\end{cases}$ (3)
with each $R_{i}\subseteq R$. Focussing on traction, this gives a total, $\boldsymbol{\mathcal{S}}_{T}$, of slip processes as
$\boldsymbol{\mathcal{S}}_{T}(x,y,\boldsymbol{u}_{s})=\sum_{i\in\{\alpha,\beta,\gamma,...\}}I_{R_{i}}(x,y)\boldsymbol{\mathcal{S}}_{R_{i}}(\boldsymbol{u}_{s})\,$ (4)
where the subscript $R_{i}$ refers to a slip process occurring in $R_{\alpha}$, $R_{\beta}$ and so on (Fig. 6) and $\boldsymbol{u}_{s}=(u_{s,x},u_{s,y})$ refers to the slip velocity at the ice-bed interface within each $R_{i}$ which can be defined using local basis vectors. For practical applications to discretised grid cells $b_{io}$, the inner-outer boundary layer, should therefore define a plane. Appendix A3 shows an example of $\boldsymbol{\mathcal{S}}_{T}$ expanded into cases for three overlapping regions. This formulation furthermore follows most work on glacier sliding in assuming velocity is continuous in time (i.e. not stick-slip). Similarly for distributed processes
$\boldsymbol{\mathcal{D}}_{T}(x,y,\boldsymbol{u},b_{b},b_{io})=\sum_{i\in\{a,b,c,...\}}I_{R_{i}}(x,y)\boldsymbol{\mathcal{D}}_{R_{i}}(\boldsymbol{u},b_{b},b_{io})\,$ (5)
where $b_{b}$ is the position of the ice-bed interface, and $\boldsymbol{u}=(u_{x},u_{y},u_{z})$ refers to the velocity field within the zone of deformation considered. Geometric resistance is then defined as
$\boldsymbol{\mathcal{G}}(F_{n},\boldsymbol{n})=-I_{F_{n}}(F_{n})\binom{F_{n}\cdot n_{x}}{F_{n}\cdot n_{y}}$ (6)
where $F_{n}$ is the normal force at a given point and $\boldsymbol{n}=(n_{x},n_{y},n_{z})$ is the corresponding unit normal vector defining the orientation of the ice-bed interface.$I_{F_{n}}(F_{n})$ is another indicator function defined as
$I_{F_{n}}(F_{n})=\begin{cases}1&\text{if }F_{n}\geq 0\,,\\ 0&\text{if }F_{n}<0\,.\end{cases}$ (7)
used in order to avoid integrating over regions of negative normal force. In practice and at longer (i.e. annual) time scales there should be no persistent negative normal force. If cavities (or subglacial lakes) are present there are two strategies to deal with them under this framework. First, we can assume the cavities are fully developed and take the water-ice interface where $F_{n}\simeq 0$ as the de facto ice-bed interface. The interface between ice and water can then simply be accommodated with a $\boldsymbol{\mathcal{S}}_{R_{i}}$ function (Eq. 4) through a frictionless relationship. Second, if $R$ covers a sufficiently large area, cavitation processes at a small scale can be subsumed within $\boldsymbol{\mathcal{S}}_{T}$ as a slip process (e.g. Helanow and others,, 2021). We are therefore not proposing a new way to simulate cavities, rather a new way to incorporate them into a larger framework.
$\boldsymbol{\mathcal{S}}_{T}$, $\boldsymbol{\mathcal{D}}_{T}$, and $\boldsymbol{\mathcal{G}}$ can then be integrated over the region $R$, with area $A_{R}$ to provide the area-averaged traction at the model bed interface
[…]
which could additionally be modified to include e.g. distributed deformation within a sediment layer through an additional term within the double integral.
Figure 6: Sub-regions related to sub-processes within a given region $R$. $R_{\alpha}$ and $R_{\beta}$ refer to slip sub-processes (Eq. 4), with the sub-region labelled $R_{\alpha}\cap R_{\beta}$ featuring both $R_{\alpha}$ and $R_{\beta}$. $R_{a}$, $R_{b}$, and $R_{c}$ refer to distributed deformation processes (Eq. 5) extend over height $h$. Dot-dash lines indicates the possible extension of region $R$ from a curved shape to a rectangular grid cell. We keep the vertical positions of $b_{b}$ and $b_{io}$ constant in this diagram, but this does not need to be the case.
It is further possible to remedy our agnosticism of the functions relating traction to velocity through the use of generic constitutive relationships for slip and deformation. Setting a given
$\boldsymbol{\mathcal{S}}_{R_{i}}=-I_{R_{i}}(x,y)P_{R_{i}}\left(\frac{|\boldsymbol{u}_{s}|}{|\boldsymbol{u}_{s}|+{u_{t}}_{R_{i}}}\right)^{\frac{1}{m_{R_{i}}}}\frac{\boldsymbol{u}_{s}}{|\boldsymbol{u}_{s}|}$ (9)
where $P_{R_{i}}$, and $m_{R_{i}}$, and ${u_{t}}_{R_{i}}$ are constants for a given region $R_{i}$ and which can represent the regularised-Coulomb relationship in Table 1, with the possibility to additionally incorporate $N$ if it is considered to be fully parametrised within $P$, or in a similar manner for a power-law relationship. A generic viscous distributed deformation sub-process can then be
[…]
where $b_{b}$ is the position of the ice-bed interface, $b_{io}$ is the height of the inner-outer flow boundary (Fig. 6), $\dot{\epsilon}_{x}^{2}=\dot{\epsilon}_{xi}\dot{\epsilon}_{xi}$ and $\dot{\epsilon}_{y}^{2}=\dot{\epsilon}_{yi}\dot{\epsilon}_{yi}$ are the effective strain rates in the $x$ and $y$ directions respectively, obtained from the strain rate tensor $\dot{\boldsymbol{\epsilon}_{ij}}=\frac{1}{2}\left(\frac{\partial u_{i}}{\partial x_{j}}+\frac{\partial u_{j}}{\partial x_{j}}\right)$. Similarly to Eq. 9, $K_{R_{i}}$ and $n_{R_{i}}$ are constants for a given region $R_{i}$, but can be related to the parameters $A$ and $n$ respectively in the commonly used Nye-Glen isotropic flow law (Nye,, 1952; Glen,, 1955), $\dot{\boldsymbol{\epsilon}}=A\tau_{e}^{n-1}\boldsymbol{\tau}$, where $\tau_{e}^{2}=\frac{1}{2}\mathrm{tr}(\boldsymbol{\tau}^{2})$ is the effective stress and $\boldsymbol{\tau}$ is the deviatoric stress tensor.
It is more challenging to suggest a boilerplate equation for $\boldsymbol{\mathcal{G}}$. Lithostatic pressure will balance, and $P_{w}$ shouldn’t have a directional influence if we continue taking the ice-water interface as the de facto ice-bed interface, so dynamic pressure and surface slope will control the net effect of $F_{n}$ integrated over $R$, as studied by e.g. Weertman, (1957), Nye, (1970). At very low Reynolds numbers low-angled obstacles provide minimal or negligible resistance, but steep and abrupt obstacles may present important resistance contributions, which cannot be calculated analytically except in special cases (Fox and others,, 2006). Usually, the resistance, $F_{D}$, from a single more prominent obstacle is defined as
$F_{D}=C_{D}\frac{1}{2}\rho V^{2}A$ (11)
where $C_{D}$ is the drag coefficient specific to that obstacle and usually obtained experimentally, $V$ is the background velocity around the obstacle, and $A$ is the cross-sectional area of the obstacle opposing flow (Fox and others,, 2006, 9.7). Considering a sphere moving through a creep-dominated fluid with very low (<<1) Reynolds number, the scaling relationship is
$F_{D}\sim K_{R_{i}}V^{\frac{1}{n}}L^{2-\frac{1}{n}}$ (12)
(Appendix A2) where $L$ is the length scale of the obstacle. Therefore, while we stop short of suggesting an explicit relationship for $\boldsymbol{\mathcal{G}}$ here, given the uncertainty in its form for potentially complicated bed geometries, it is sufficient to say for now that $\boldsymbol{\mathcal{G}}$ will have negligible importance for very smooth beds, but be an increasing function of velocity, obstacle prominence, $\mathcal{O}_{p}$, and obstacle frequency, $\mathcal{O}_{p}$, and which may follow a power law similar to $\boldsymbol{\mathcal{D}}_{R_{i}}$. We represented this functionally as
$\boldsymbol{\mathcal{G}}=f(\boldsymbol{u}_{b},\mathcal{O}_{p},\mathcal{O}_{f})\,.$ (13)
We leave it to future studies to better quantify $\boldsymbol{\mathcal{G}}$ in the context of realistically rough landscapes, but provide Eq. 13 as scaffolding in the meantime.
Overall, this framework extends from Fowler1977 to include arbitrarily many sub-processes in three dimensions and facilitates quantification of the role of each sub-process in the inner flow, though a full representation will be further complicated in reality; for example if we consider multiple rheological layers between $z=b_{b}$ and $z=b_{io}$. Form drag is covered in greater detail later (Section 3.3), but we henceforth intend the sum of $\boldsymbol{\mathcal{D}}$ and $\boldsymbol{\mathcal{G}}$ when we refer to form drag processes. We consider how this framework could relate to a unified sliding relationship in Section 6.3.
@@ 3.1 Sliding over ‘soft’ (sediment) beds
More recently, Zoet and Iverson, (2020) used a ring of temperate ice $\sim$15 cm thick between the rigid shear ring and underlying sediment. In this setting, rate-strengthening behaviour occurs at lower velocities (below around 50 m a-1 in sediment with clasts, Fig. 4a) as ice viscously deforms and regelates around static clasts before the clasts begin to plough through the finer-grained matrix at a stress limited by the till’s Coulomb strength (Fig. 4b). The net result of the Zoet and Iverson, (2020) experiment is a regularised-Coulomb relationship (Fig. 1, Table 1) where the sliding interface is defined up 15 cm above the ice-till interface.
@@ 3.2 Sliding over ‘hard’ (bedrock) beds
The slip of ice over a planar surface (Fig. 4c) is closer to the assumptions of standard Coulomb behaviour (Coulomb,, 1785; Desplanques,, 2014), though plane-normal velocity components, englacial clasts, and possible clast-scale cavitation present complications (Hoffman and others,, 2022). Rather than a sliding relationship of the form of Eq. 1, basal traction in these studies is often reported as a static $\mu_{s}$ or transient $\mu_{t}$ coefficient of friction, $\mu_{s}\>\textup{or}\>\mu_{t}=\frac{\tau_{b}}{N}$ where $\mu_{s}$ describes the force required to move the plane from stationary and $\mu_{t}$ describes the force required to keep the plane in motion. Shear-ring studies using temperate ice and englacial sediment content of 0-20% over macroscopically flat granite find a $\mu_{s}$ of 0.02-0.05 (Barnes and others,, 1971; Zoet and others,, 2013; McCarthy and others,, 2017; Thompson,, 2020), lower than teflon-on-teflon with no lubrication (Fetfatsidis and others,, 2013). In the above studies, sliding may be stick-slip or steady-state (outlined further in Section 3.5) depending on water pressure and particular material properties. Transient coefficients of friction between 0.05-0.08 were obtained from subglacial access to the ice-rock interface of the Engabreen outlet glacier in Norway (Iverson and others,, 2003; Cohen and others,, 2005). Unexpectedly high shear traction values of up to 0.5 MPa accompany the low $\mu_{t}$ values at Engabreen and these are still not fully explained, suggesting a possible gap in understanding – see Thompson and others, (2020) for a more in depth discussion (and note that, in calculations of transient friction, in contrast to static friction, the line defined by $\mu_{t}$ will not necessarily pass through the graph origin of $N$ and $\tau_{b}$). The transient coefficient of friction is much higher for cold ice (0.5 at -20oC) making sub-temperate slip very low, but potentially non-negligible (Barnes and others,, 1971; McCarthy and others,, 2017; Atkins,, 2013; Mantelli and others,, 2019; Section 3.7).
@@
Sliding over rough hard beds has also received much attention. This began with the classical work of Weertman, (1957), who considered ice motion over isolated cuboids (akin to the geometry in Fig. 4e) with no ice-bed separation through two processes: regelation and enhanced creep (Fig. 4e). In regelation, ice moves without deformation. Temperate ice approaching an obstacle is subject to greater pressure, lowering its melting point. Consequently, some of the ice melts at the upstream face, and the meltwater is driven to the downstream face by the pressure gradient. Here, where the pressure is lower and the melting point is higher, the meltwater refreezes. The higher temperature at the downflow side of the obstacle results in heat flow back through the obstacle, completing the cycle. Regelation is an ongoing process in many subglacial settings (Kamb and Lachapelle,, 1964; Kamb,, 1970; Hallet and others,, 1978; Hubbard and Sharp,, 1993; Iverson,, 2000; Cook and others,, 2011; Rempel and Meyer,, 2019), including in soft bed settings outside the scope of the original Weertman, (1957) paper. However, the requirement for return heat flow causes regelation rate to vary inversely with obstacle wavelength, limiting the importance of regelation as a sliding process to obstacles smaller than a few cm (Weertman,, 1957). Regelation is therefore generally not viewed as a significant contributor to overall sliding rates (MacAyeal,, 2019) and has received comparatively little attention since the mid 1990s, though further contemporary study would not be unwelcome.
@@
Enhanced creep is prompted by stress concentrations at bedrock obstacles, leading to locally increased strain rates. Functionally, this means enhanced creep can be considered as a synonym for form drag (Section 3.3), and would fall under $\boldsymbol{\mathcal{D}}_{T}$ (Eq. 5) in our framework (Section 2). Conversely to regelation, the contribution of enhanced creep to glacier sliding in Weertman’s 1957 study increases with obstacle size. Focusing on the obstacle sizes that produce the greatest rate of sliding when regelation and enhanced creep act in unison (millimetre- to metre-scale) Weertman, (1957) reaches a value of $m$ in a power-law of $(n+1)/2=2$ with $C$ representing a combination of bed-geometry characteristics. Technically, values of $m$ deviating from 2 when $n=3$ are therefore not physically based on the original Weertman theory making ‘power-law’, which we use henceforth, or ‘Weertman-type’ sliding a more suitable descriptor. Sliding over a sinusoidal two-dimensional bed with no ice-bed separation and a linear ice rheology produces a linear sliding relationship ($m=1$ in power-law sliding) (Nye,, 1969), as does Weertman sliding with $n=1$.
@@
power-law sliding, usually with $m=3$, is the most commonly used sliding relationship in the latest round of ISMIP experiments (with linear sliding also frequently used) (Goelzer and others,, 2020; Seroussi and others,, 2020), yet the theoretical underpinnings supporting these relationships in the original papers do not hold up to scrutiny – including by Weertman’s own reappraisal (Weertman,, 1979). Weertman’s original theory features a non-linear rheology following Glen, (1955), but it is apparent that no glacier beds are actually characterised by a controlling obstacle size with the morphology used by Weertman, (1957), and even changes to a different unrealistic bed geometry produces a materially different relationship (Lliboutry,, 1979). Weertman sliding is further challenged by ubiquitous evidence and theoretical support for subglacial cavitation, or regions of bounded ponded water at high pressure between the glacier sole and underlying hard or soft bed, which form when the pressure on the downstream face of an obstacle falls below a critical level (Fig. 4g and e.g. Walder and Hallet,, 1979; Kamb,, 1987; Hooke,, 1989; Helanow and others,, 2021). Weertman sliding also does not provide an explanation for increased sliding rates when subglacial water pressure increases. The persistent utility of power-law sliding in heuristic studies, though generally with $m\geq 3$, despite these major problems is covered in Sections 5 and 6.3.
@@
Subglacial cavities (Fig. 4g) act to reduce contact between the glacier sole and the ice bed and may grow as a result of increasing basal velocity or subglacial water pressure. In the two-dimensional formulations of Lliboutry, (1968), Iken, (1981), Schoof, (2005), and Gagliardini and others, (2007), amongst others, where the bed is comprised of frictionless sinusoids, this produces an upper bound for basal traction, followed by rate-weakening behaviour (Fig. 1). Put mathematically, ‘Iken’s bound’ as coined by Schoof, (2005) is given as
@@
where $\theta$ is the maximum up-slope angle between the bed and the mean flow direction of the ice (Iken,, 1981; Schoof,, 2005). Roldan-Blasco and others, (2022) suggest that in the simplified two-dimensional framework of compound sinusoids non-negligible friction at the ice-bed interface will not significantly alter the form of the sliding relationship or the existence of Iken’s bound. However, the geometry used in these studies can also be considered unrealistic, with the absence of a third dimension preventing lateral escape of subglacial water which would (1) act to reduce the cavity size (Gimbert and others,, 2021) and (2) may place a limit on the maximum obtainable subglacial water pressure averaged over a large region at long (i.e. annual) time scales after the system has had time to equilibrate (e.g. Dow and others,, 2015; De Fleurian and others,, 2018; Hart and others,, 2022). Helanow and others, (2019) and Helanow and others, (2021) address the three-dimensional problem for smaller scales, and show that realistic three-dimensional bed topographies up to 25 m still result in bounded basal traction, but that significant rate-weakening behaviour is not expected (the regularised-Coulomb relationship in Fig. 1). The validity of the assumptions underlying Eq. 14 when considering inner flows over larger areas and rough topography is covered in Section 6.2.